ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 197, 2013

Alesker, S. and E. Shelukhin On a uniform estimate for the quaternionic Calabi problem

309

Altınel, T., J. Burdges and O. Frécon On Weyl groups in minimal simple groups of finite Morley rank

377

Alves, M. M. S., E. Batista, M. Dokuchaev and A. Paques Twisted partial actions of Hopf algebras

263

Barshevsky, O., M. Fayers and M. Schaps A non-recursive criterion for weights of a highest-weight module for an affine Lie algebra

237

Batista, E. See M. M. S. Alves Bialy, M. Smooth solutions for a p-system of mixed elliptic-hyperbolic

type

189

Blagojević, P. V. M., B. Bukh and R. Karasev Turán numbers for K_{s,t}-free graphs: topological obstructions and algebraic constructions

199

Bourgain, J. On the Fourier–Walsh spectrum of the Moebius function

215

Bowman, J. P. and F. Valdez Wild singularities of flat surfaces  69

Brooke-Taylor, A. D. and S.-D. Friedman Subcompact cardinals, squares, and stationary reflection

453

Bukh, B. See P. V. M. Blagojević Burdges, J. See T. Altınel Chapman, A. and U. Vishne Square-central elements and standard

generators for biquaternion algebras

409

Chuang, C.-L. Automorphisms of Ore extensions  437
Collins, M. J. Induced representations and invariant forms  51
Cox, S. and M. Viale Martin’s Maximum and tower forcing  347

Dokuchaev, M. See M. M. S. Alves Fayers, M. See O. Barshevsky Frécon, O. See T. Altınel Friedland, O. and O. Guédon Sparsity and non-Euclidean embed-

dings

329

Friedman, S.-D. See A. D. Brooke-Taylor González-Aguilera, V. and A. Liendo On the order of an auto-

morphism of a smooth hypersurface 29 Green, D. J. and J. Lynd Weak closure and Oliver’s p-group con-

jecture

497

Guédon, O. See O. Friedland Heckenberger, I. and H.-J. Schneider Right coideal subalgebras

of Nichols algebras and the Duflo order on the Weyl groupoid

139

Isaacs, I. M., M. Loukaki and A. Moretó The average degree of an irreducible character of a finite group

55

Janko, Z. Finite p-groups all of whose cyclic subgroups A, B with A ∩ B ≠ {1} generate an abelian group

1

Karasev, R. See P. V. M. Blagojević Kim, S. K. The Bishop–Phelps–Bollobás Theorem for operators from

c_{0} to uniformly convex spaces

425

Li, S. and A. Naor Discretization and affine approximation in high dimensions

107

Liendo, A. See V. González-Aguilera Loukaki, M. See I. M. Isaacs Lynd, J. See D. J. Green Moretó, A. See I. M. Isaacs Naor, A. See S. Li Nayar, P. and T. Tkocz The unconditional case of the complex

S-inequality

99

Nuño-Ballesteros, J. J., B. Oréfice-Okamoto and J. N. Tomazella The vanishing Euler characteristic of an isolated determinantal singularity

475

Oréfice-Okamoto, B. See J. J. Nuño-Ballesteros Paques, A. See M. M. S. Alves Petrov, F. and P. Zusmanovich On Shirshov bases of graded al-

gebras

23

Prelli, L. De Rham theorem for Schwartz functions on Nash manifolds

131

Schaps, M. See O. Barshevsky Schneider, H.-J. See I. Heckenberger Shelukhin, E. See S. Alesker Tkocz, T. See P. Nayar Tomazella, J. N. See J. J. Nuño-Ballesteros Valdez, F. See J. P. Bowman Viale, M. See S. Cox Vishne, U. See A. Chapman Zusmanovich, P. See F. Petrov